A vessel of volume 5 liters was filled with 2 moles of hydrogen gas (a), then a further





1 mole of chlorine gas (b), and then the mixture was sparked leading to the formation of 1 mole of HCI (g) (c). What are the partial pressure of the components





and the total pressure of the system at the final stage. The temperature being 25°C throughout.

1
Expert's answer
2022-02-18T09:56:02-0500

According to the reaction

H2 + Cl2 --> 2HCl,

if 1 mole of HCl was formed, 0.5 moles of each H2 and Cl2 have actually reacted. Therefore, the final amounts of each gas are:

H2: 2 - 0.5 = 1.5 mol

Cl2: 1 - 0.5 = 0.5 mol

HCl: 1 mol

Now, partial pressures can be determined:


"P_{H_2}=\frac{n_{H_2}RT}{V}=\frac{1.5\ mol\times0.08206\frac{L\cdot{atm}}{mol\cdot{K}}\times298\ K}{5\ L}=7.34\ atm"


"P_{Cl_2}=\frac{n_{Cl_2}RT}{V}=\frac{0.5\ mol\times0.08206\frac{L\cdot{atm}}{mol\cdot{K}}\times298\ K}{5\ L}=2.45\ atm"


"P_{HCl}=\frac{n_{HCl}RT}{V}=\frac{1\ mol\times0.08206\frac{L\cdot{atm}}{mol\cdot{K}}\times298\ K}{5\ L}=4.89\ atm"


According to the Dalton's law,

"P_{total}=P_{H_2}+P_{Cl_2}+P_{HCl}=7.34+2.45+4.89=14.7\ atm"


Answer: H2 - 7.34 atm; Cl2 - 2.45 atm; HCl - 4.89 atm; total - 14.7 atm